# Segmentation-free Goodness of Pronunciation

### v2

@@ VI Conclusions
In this work, we propose improvements to a framework for pronunciation assessment that we recently introduced in [cao24b_interspeech]. We propose two methods with the goal of allowing the use of goodness of pronunciation (GOP) features extracted from modern high-performance automatic speech recognition models. The main idea of our methods is to release the assumption that the GOP evaluation should be performed on a predetermined segment of speech corresponding to the target phone. We do this either by adjusting the segment under test to the activations of the acoustic models (self-aligned GOP, GOP-SA) or by proposing a definition of GOP that is independent of the target segment alignment (alignment-free GOP, GOP-AF).
We provide theoretical derivations that make the underlying assumptions for GOP-AF explicit. We also provide experimental results exploring a number of properties of our methods. We show how GOP-SA is always better than traditional GOP regardless of the characteristics of the acustic model used for the evaluation. We show that GOP-AF obtains the overall best results both for the CMU-Kids and for the SpeechOcean762 material. On the SpeechOcean762 data we obtain state-of-the-art results keeping the MDD model constant. We also show how the peakiness of the acoustic models affect the MDD results for the standard GOP definition and for the two proposed methods. Finally we show that the performance of GOP-AF, that considers the full utterance to assess the pronunciation of the target phone, is not affected by the length of the context.

### v3

@@ VI Conclusions
In this work, we propose improvements to a framework for pronunciation assessment that we recently introduced in [4]. We propose two methods with the goal of allowing the use of goodness of pronunciation (GOP) features extracted from modern high-performance automatic speech recognition models. The main idea of our methods is to release the assumption that the GOP evaluation should be performed on a predetermined segment of speech corresponding to the target phone. We do this either by adjusting the segment under test to the activations of the acoustic models (self-aligned GOP, GOP-SA) or by proposing a definition of GOP that is independent of the segmentation of test utterance (segmentation-free GOP, GOP-SF).
Our theoretical derivations make the underlying assumptions for GOP-SF explicit. We also provide experimental results exploring a number of properties of our methods. GOP-SA is always better than traditional GOP regardless of the characteristics of the acoustic model used for the evaluation. GOP-SF obtains the overall best results for both the CMU Kids and for the speechocean762 material. On the speechocean762 data we obtain state-of-the-art results keeping the MDD model constant. We also show how the peakiness of the acoustic models affects the MDD results for the standard GOP definition and for the two proposed methods. Finally we show that the performance of GOP-SF, that considers the full utterance to assess the pronunciation of the target phone, is not affected by the length of the context.